### v1

@@ 6.1 State Features Collaborate to Determine Value Estimates
We now take a game-theoretic view of value estimation. The estimated value $\hat{v}^{\pi}(s)$ can also be viewed as the outcome of a collaborative process among features. That is, each feature value in $s$ contributes to the estimation of return under policy $\pi$. To attribute this estimate across the feature values, we define another coalitional game over the features, introduce a new set of fairness axioms tailored to the estimation setting, and identify the Shapley values as the unique solution.
Definition 16 (Value Estimation Game)
A set =${\cal F}=\{1,\dots,n\}$ of features and a characteristic function $\hat{v}^{\pi}_{s}:2^{|{\cal F}|}\rightarrow\mathbb{R}$, where $\hat{v}^{\pi}_{s}(\mathcal{C})$ denotes the estimated expected return when only the feature values $s_{\mathcal{C}}\in{\cal S}_{\mathcal{C}}$ are observed.
As before, we postpone the precise construction of the characteristic function $\hat{v}^{\pi}_{s}$ until later. For now, we ask: what is the contribution of each feature value $s_{i}\in{\cal S}_{i}$ to the estimate $\hat{v}^{\pi}(s)$? We denote this contribution by $\phi_{i}(\hat{v}^{\pi}_{s})$, and define four axioms that any such attribution should satisfy.
  • Axiom 1 (Efficiency). The difference in value between the estimated return when all features are known, $\hat{v}^{\pi}_{s}({\cal F})$, and when none are known, $\hat{v}^{\pi}_{s}(\emptyset)$, is fully distributed across the features:
    $\hat{v}^{\pi}_{s}({\cal F})-\hat{v}^{\pi}_{s}(\emptyset)=\sum_{i\in{\cal F}}\phi_{i}(\hat{v}^{\pi}_{s})$
  • Axiom 2 (Linearity). For a composite policy $\alpha\,\pi_{1}+(1-\alpha)\,\pi_{2}$, a feature’s contribution to the estimated expected return under the combined policy is equal to the same weighted sum of its contributions under each individual policy:
    $\phi_{i}(\alpha V^{\pi_{1}}_{s}+(1-\alpha)V^{\pi_{2}}_{s})=\alpha\,\phi_{i}(V^{\pi_{1}}_{s})+(1-\alpha)\,\phi_{i}(V^{\pi_{2}}_{s})$
  • Axiom 3 (Symmetry). Two features receive equal attribution if they make identical marginal contributions to the estimated return for all possible coalitions:
    $\phi_{i}(\hat{v}^{\pi}_{s})=\phi_{j}(\hat{v}^{\pi}_{s})\quad\text{if}\quad\hat{v}^{\pi}_{s}(\mathcal{C}\cup\{i\})=\hat{v}^{\pi}_{s}(\mathcal{C}\cup\{j\})\quad\forall\mathcal{C}\subseteq{\cal F}\setminus\{i,j\}$
  • Axiom 4 (Nullity). A feature receives zero attribution if it makes no marginal contribution to the estimated return in any coalition:
    $\phi_{i}(\hat{v}^{\pi}_{s})=0\quad\text{if}\quad\hat{v}^{\pi}_{s}(\mathcal{C}\cup\{i\})=\hat{v}^{\pi}_{s}(\mathcal{C})\quad\forall\mathcal{C}\subseteq{\cal F}\setminus\{i\}$
These four axioms mirror those used when explaining behaviour and performance.
@@
The Shapley values $\phi_{1}(\hat{v}^{\pi}_{s}),\dots,\phi_{n}(\hat{v}^{\pi}_{s})$ for the value estimation game $({\cal F},\hat{v}^{\pi}_{s})$ are the unique allocation of the estimated return $\hat{v}^{\pi}(s)$ across the features of state $s$ that satisfies the four axioms of fair credit assignment: efficiency, linearity, symmetry, and nullity.
The Shapley values of the value estimation game provide a principled way to attribute state-value estimates to individual state features, offering a structured explanation of value estimation. As with the previous coalitional games, computing these Shapley values requires precisely defining the characteristic function $\hat{v}^{\pi}_{s}(\mathcal{C})$, which we turn to next.
@@ 6.2 What Can Be Assumed About Value Estimates When Some State Features Are Unknown?
We now consider how to define the characteristic function $\hat{v}^{\pi}_{s}(\mathcal{C})$: the estimated return from state $s$ when only the features in $\mathcal{C}\subseteq{\cal F}$ are observed. A natural principle is to estimate value using only the information that is actually available. The estimated return should reflect the best assessment based on the observed features alone.
@@
Let $\hat{u}^{\pi}:{\cal S}_{\mathcal{C}}\rightarrow\mathbb{R}$ be a value function defined over a subset of features $\mathcal{C}\subseteq{\cal F}$ such that the estimated return from state $s$ when only the features in $\mathcal{C}$ are observed is given by:
@@
When comparing return estimates using mean squared error, the function $\hat{u}^{\pi}$ is the closest in expectation to $\hat{v}^{\pi}$ among all functions defined over the subset of features $\mathcal{C}$:
@@
This motivates our definition of the characteristic function for value estimation:
Definition 20 (Value Estimation Characteristic Function)
The estimated return from state $s$ when only the features in $\mathcal{C}$ are observed is given by:
@@
When only the features in $\mathcal{C}$ are observed, the function $\hat{u}^{\pi}$ provides the best approximation of the full-information value estimate $\hat{v}^{\pi}(s)$. As such, the difference between $\hat{u}^{\pi}(s_{\mathcal{C}})$ and $\hat{v}^{\pi}(s)$ quantifies how necessary the unobserved features $\bar{\mathcal{C}}$ are for recovering the original estimate. Defining the characteristic function as $\hat{v}^{\pi}_{s}(\mathcal{C})=\hat{u}^{\pi}(s_{\mathcal{C}})$ ensures that the differences across feature subsets used to compute Shapley values reflect the influence of the unknown features, making it a principled choice for the value estimation game.
As with explaining policy and performance, this construction is general to any policy. Alternative characteristic functions could also be used. In particular, a marginal variant of $\hat{v}^{\pi}_{s}(\mathcal{C})$—which assumes feature independence—may be suitable in some settings. An analogous discussion of this alternative appears for the policy characteristic function in Section 4.3.
@@ 6.3 Illustrative Example
  • $(\text{R},10)$: The initial state with a right-facing sign and distance 10 miles.
  • $(\text{L},2)$: The second state, reached by turning right, with a left-facing sign and distance 2 miles.
Using the characteristic function introduced above, we now apply our framework to analyse how value is estimated in each state. Table 6 presents the return estimates for different feature subsets, and the resulting Shapley values for each feature.
Table 6: The steady-state distribution, characteristic values, and Shapley values when explaining value estimation for each state in the road sign example (Figure 1).
@@
We start by focusing on the contribution of the Distance feature. In state $(\text{R},10)$, the agent is 10 miles from the goal. The distance feature reduces the return estimate relative to the baseline with no features, resulting in a negative Shapley value ($-0.25$). In state $(\text{L},2)$, the agent is much closer to the goal, and the distance feature increases the return estimate, producing a positive Shapley value ($+0.25$).
These attributions reflect an intuitive interpretation: when the agent is far from the goal, the expected return is lower; when it is near, the expected return is higher. The Shapley values capture how each feature contributes to the estimate, correctly assigning responsibility to the Distance feature in line with how it influences return estimation.
The Direction feature receives identical attributions because it also helps differentiate the two visited states. In this domain, the sign’s direction indirectly conveys distance: the policy only encounters the right-facing sign when far from the goal and the left-facing sign when close.
This example demonstrates how our framework enables Shapley values to provide meaningful explanations of value estimation, revealing how individual features guide our estimate of the agent’s expected outcomes. [^3] See Section B.3 for an additional example illustrating how SVERL explains value estimation.
@@ 7 Understanding SVERL Through Examples
We present two examples using SVERL to explain agent-environment interactions in practice. The first example shows how explanations of behaviour, performance, and value estimation can align or diverge to reveal complementary insights. The second example applies the framework in a larger domain, demonstrating its scalability to the feasible limits of exact Shapley value computation. Unless stated otherwise, all examples are undiscounted and explain optimal policies computed via Value Iteration (Sutton and Barto, 2018). Steady-state distributions are either computed exactly or accurately estimated from trajectories; expected returns for the performance characteristic are learnt using Value Iteration. [^4] Code is available at https://github.com/djeb20/sverl.
@@ 7.1 Mastermind
Figure 2: Shapley value attributions for three successive states in Mastermind (top to bottom). First column: episode progression. Second column: feature contributions to the probability of selecting the optimal action taken (behaviour). Third column: contributions to the agent’s performance. Fourth column: contributions to the agent’s return estimate (value estimation). Shapley values are scaled to lie between $-1$ and $1$.
The first domain we consider is Mastermind, based on the eponymous code-breaking board game. In Mastermind, at the start of each episode, the environment randomly samples one of four codes: $\mathrm{AA}$, $\mathrm{AB}$, $\mathrm{BA}$, or $\mathrm{BB}$, hidden from the agent. At each decision stage, the agent presents one of the codes. The environment then returns two types of feedback to indicate how well the agent’s code matches the hidden code:
@@
  • Position clue: the number of letters that are both correct and in the correct position.
@@
  • Misplaced clue: the number of correct letters that are in the wrong position.
These two clues are calculated independently, and a letter contributing to the position clue cannot also contribute to the misplaced clue. Incorrect codes incur a reward of $-1$ and correct codes incur a reward of 00. The game terminates at the end of three decision stages, unless the agent identifies the code correctly in an earlier decision stage. Each state is represented as a game board, with up to three agent codes and each corresponding pair of clues. Each agent code has four associated features: two for the code and two for the feedback it received. If fewer than three codes have been presented, the remaining features take on an empty value.
We use SVERL to analyse an example episode where the hidden code is $\mathrm{AB}$. Figure 2 presents the trajectory for the agent’s policy: deterministically selecting $\mathrm{AA}$, $\mathrm{BA}$, then $\mathrm{AB}$, and explanations of behaviour, performance, and value estimation.
We start with the initial decision stage, where the agent has not yet presented a code. In other words, the board is clear. The agent’s policy deterministically selects the action $\mathrm{AA}$. The explanation of behaviour highlights the empty bottom row—corresponding to the agent’s first code—as the largest contributor towards the probability of this action. This explanation reflects that the agent’s policy chooses $\mathrm{AA}$ precisely because it’s in the initial state where no prior information is available. Similarly, the explanation of value estimation highlights this row too, but for decreasing the agent’s estimated return. Again, this reflects that the empty bottom row indicates the environment is in the first state, with the lowest expected return. In contrast, the performance explanation attributes no responsibility to any feature, correctly indicating that all possible initial agent codes result in maximum expected return.
In the second state, the first code $\mathrm{AA}$ and its clues have revealed: one of the $\mathrm{A}$s is in the correct position, and the other should be a $\mathrm{B}$. The agent now selects $\mathrm{BA}$. The behaviour and performance explanations identify the position clue as the primary driver of action selection and return. While the misplaced clue is informative, it is 00 for the initial code $\mathrm{AA}$ with any code, and thus does not meaningfully influence the agent’s decisions. For value estimation, the position clue of $1$ receives a negative attribution: a value of 00 would have revealed the hidden code and yielded a higher expected return, so the observed value of $1$ lowers the estimate.
@@
In the final state, the agent correctly proposes $\mathrm{AB}$. The previous code and clues uniquely determine the hidden code and the correct action. The features corresponding to the code and clues receive high attribution across all three explanations: they reveal the optimal move (behaviour), justify the return received (performance), and identify a state with maximal return (value estimation).
This example illustrates how SVERL captures the evolving role of information in decision making, outcomes, and predictions as the game progresses.
@@ 7.2 Minesweeper
Figure 3: Shapley value attributions for two successive states in Minesweeper (top to bottom). First column: episode progression. Second column: feature contributions to the probability of selecting the optimal action taken (behaviour). Third column: contributions to the agent’s performance. Fourth column: contributions to the agent’s return estimate (value estimation). Mines are hidden from the agent; markers “M1” and “M2” are shown for reference. Shapley values are scaled to lie between $-1$ and $1$.
Minesweeper is played on a $4\times 4$ grid with two hidden mines randomly placed at the start of each episode and all squares unopened. The agent selects unopened squares to reveal either a mine-free square with a number indicating how many mines total are in adjacent squares, or a mine, which ends the episode with a reward of $-1$. If an opened square shows 00, adjacent squares are opened recursively until reaching nonzero borders. If all safe squares are revealed, the episode ends with a reward of 00. This domain presents a more complex and higher-dimensional challenge. Each state is represented as a game board, consisting of 16 features, each taking one of the following values: the number of adjacent mines (0, 1, or 2) or the value unopened. The domain has approximately $175,000$ states. We explain the behaviour, performance, and value estimation of a policy learnt via Q-learning (Watkins and Dayan, 1992), with the performance characteristic estimated using Monte Carlo rollouts. We consider the two states shown in Figure 3.
In the first state, the optimal action is to open square $(4,2)$. Many features provide information about the first mine (M1), but only the clue under square $(4,2)$ identifies the location of the second mine (M2). The clue is thus pivotal to safe and optimal play. The behaviour and performance explanations both assign high attribution to this feature. In Minesweeper, states with more unopened squares carry more uncertainty about mine locations and, therefore, have lower expected returns. This is reflected by the explanation of value estimation, where unopened squares receive negative attributions, and opened squares receive positive attributions.
In the second state, the agent opens square $(4,4)$. The clue revealed at $(4,2)$ by the previous action has uniquely identified both mine locations, determining the optimal action and return. The feature corresponding to the clue again receives high attribution in the behaviour and performance explanations. As before, opened squares contribute positively and unopened squares negatively to value estimation, reflecting how greater certainty leads to higher return estimates.
This example shows how our framework scales to more complex environments, revealing which features drive safe decision-making and how uncertainty affects expected outcomes.

### v2

@@ 6.1 State Features Collaborate to Determine Value Estimates
We now take a game-theoretic view of predictions. The predicted value $\hat{v}^{\pi}(s)$ can be viewed as the outcome of a collaborative process among features. That is, each feature value in $s$ contributes to the prediction of expected return under policy $\pi$. To attribute this prediction across the feature values, we define another coalitional game over the features, introduce a new set of fairness axioms tailored to the prediction setting, and identify the Shapley values as the unique solution.
Definition 36 (Prediction Game)
A set =${\cal F}=\{1,\dots,n\}$ of features and a characteristic function $\hat{v}^{\pi}_{s}:2^{|{\cal F}|}\rightarrow\mathbb{R}$, where $\hat{v}^{\pi}_{s}(\mathcal{C})$ denotes the predicted expected return using only the feature values $s_{\mathcal{C}}\in{\cal S}_{\mathcal{C}}$.
As before, we postpone the precise construction of the characteristic function $\hat{v}^{\pi}_{s}$ until later. For now, we ask: what is the contribution of each feature value $s_{i}\in{\cal S}_{i}$ to the estimate $\hat{v}^{\pi}(s)$? We denote this contribution by $\phi_{i}(\hat{v}^{\pi}_{s})$ and appeal to Shapley’s foundational axioms, restated here to clarify their role in this specific attribution setting.
Axiom 37 (Efficiency)
The difference in value between the predicted expected return using all features, $\hat{v}^{\pi}_{s}({\cal F})$, and when using none, $\hat{v}^{\pi}_{s}(\emptyset)$, is fully distributed across the features:
$\hat{v}^{\pi}_{s}({\cal F})-\hat{v}^{\pi}_{s}(\emptyset)=\sum_{i\in{\cal F}}\phi_{i}(\hat{v}^{\pi}_{s}).$
Axiom 38 (Linearity)
For two independent policies $\pi$ and $\omega$, a feature’s contribution to the weighted sum of their predicted expected returns is equal to the corresponding weighted sum of the feature’s contributions under each policy:
$\phi_{i}(\alpha\,\hat{v}^{\pi}_{s}+\beta\,\hat{v}^{\omega}_{s})=\alpha\,\phi_{i}(\hat{v}^{\pi}_{s})+\beta\,\phi_{i}(\hat{v}^{\omega}_{s}).$
Axiom 39 (Symmetry)
Two features receive equal attribution if they make identical marginal contributions to the predicted expected return for all possible coalitions:
$\phi_{i}(\hat{v}^{\pi}_{s})=\phi_{j}(\hat{v}^{\pi}_{s})\quad\text{if}\quad\hat{v}^{\pi}_{s}(\mathcal{C}\cup\{i\})=\hat{v}^{\pi}_{s}(\mathcal{C}\cup\{j\})\quad\forall\mathcal{C}\subseteq{\cal F}\setminus\{i,j\}.$
Axiom 40 (Nullity)
A feature receives zero attribution if it makes no marginal contribution to the predicted expected return in any coalition:
$\phi_{i}(\hat{v}^{\pi}_{s})=0\quad\text{if}\quad\hat{v}^{\pi}_{s}(\mathcal{C}\cup\{i\})=\hat{v}^{\pi}_{s}(\mathcal{C})\quad\forall\mathcal{C}\subseteq{\cal F}\setminus\{i\}.$
These four axioms mirror those used when explaining behaviour and outcomes.
@@
The Shapley values $\phi_{1}(\hat{v}^{\pi}_{s}),\dots,\phi_{n}(\hat{v}^{\pi}_{s})$ for the prediction game $({\cal F},\hat{v}^{\pi}_{s})$ are the unique allocation of the predicted expected return $\hat{v}^{\pi}(s)$ across the features of state $s$ that satisfies the four axioms of fair credit assignment: efficiency, linearity, symmetry, and nullity.
The Shapley values of the prediction game provide a principled way to attribute state-value estimates to individual state features, offering a structured explanation of predictions. As with the previous coalitional games, computing these Shapley values requires precisely defining the characteristic function $\hat{v}^{\pi}_{s}(\mathcal{C})$, which we turn to next.
@@ 6.2 What Can Be Assumed About Value Estimates When Some State Features Are Unknown?
We now consider how to define the characteristic function $\hat{v}^{\pi}_{s}(\mathcal{C})$: the predicted expected return from state $s$ when using only the features in $\mathcal{C}\subseteq{\cal F}$. A natural principle is to best estimate value using only the information that is actually available. The predicted expected return should reflect the best assessment based on the observed features alone.
@@
Let $\hat{u}^{\pi}:{\cal S}_{\mathcal{C}}\rightarrow\mathbb{R}$ be a value function defined over a subset of features $\mathcal{C}\subseteq{\cal F}$ such that the estimated expected return from state $s$ when only the features in $\mathcal{C}$ are observed is given by:
@@
When comparing expected-return estimates using mean squared error, the function $\hat{u}^{\pi}$ is the closest in expectation to $\hat{v}^{\pi}$ among all functions defined over the subset of features $\mathcal{C}$:
@@
This motivates our definition of the characteristic function for predictions:
Definition 44 (Prediction Characteristic Function)
The predicted expected return from state $s$ when using only the features in $\mathcal{C}$ is given by:
@@
When only the features in $\mathcal{C}$ are observed, the function $\hat{u}^{\pi}$ provides the best approximation of the full-information value estimate $\hat{v}^{\pi}(s)$. As such, the difference between $\hat{u}^{\pi}(s_{\mathcal{C}})$ and $\hat{v}^{\pi}(s)$ quantifies how necessary the unobserved features $\bar{\mathcal{C}}$ are for recovering the original estimate. Defining the characteristic function as $\hat{v}^{\pi}_{s}(\mathcal{C})=\hat{u}^{\pi}(s_{\mathcal{C}})$ ensures that the differences across feature subsets used to compute Shapley values reflect the influence of the unknown features, making it a principled choice for the prediction game.
This construction is general to any policy. Alternative characteristic functions could also be used. In particular, a marginal variant of $\hat{v}^{\pi}_{s}(\mathcal{C})$—which assumes feature independence—may be suitable in some settings. An analogous discussion of this alternative appears for the discrete policy characteristic function in Section 4.2.
@@ 6.3 Illustrative Example
[…]
Table 6: The steady-state distribution, characteristic values, and Shapley values when explaining predictions for each state in the road sign example of Figure 1.
[…]
6.4 Contrasting Predictions with Outcomes
With the final component of SVERL introduced, we now ground the distinction between explaining outcomes and predictions in a concrete example using the game of Tic-Tac-Toe. Consider an agent playing optimally against a minimax opponent (Polak, 1989). Against such an opponent, the agent cannot win but can avoid losing. In other words, every game ends in a draw. As a result, the expected return is zero from every state. Now consider the state shown in Figure 2. The opponent (O) moved first, and it is now the agent’s (X) turn. Unless the agent marks the bottom-centre square, it will lose. We consider how each of the nine grid squares contributes to outcomes and predictions.
Figure 2: Tic-Tac-Toe. Left: Contributions to the agent’s expected return (outcome). Right: Contributions to the agent’s estimated expected return (prediction).
The left side of Figure 2 shows the SVERL attributions for explaining outcomes—how each of the nine grid squares, taking values “X”, “O”, or “empty”, contributes to the agent’s expected return through their influence on the agent’s policy. The opponent’s two squares have the highest attributions, reflecting the agent’s need to block the opponent’s winning move. The right side of Figure 2 shows the corresponding SVERL attributions for predicting the agent’s expected return. All feature contributions are zero. While initially surprising, this is correct: every state leads to a draw, so the expected return is always zero, and thus no individual feature is more or less predictive of the expected return than any other.
Explaining predictions can reveal features useful for predicting expected return, but it does not explain how feature values shape the agent’s behaviour or the outcomes of that behaviour. In contrast, explaining outcomes reveals how feature values impact expected return as a direct consequence of their impact on policy.
@@ 7 Understanding SVERL Through Examples
We present two examples to illustrate how SVERL captures the changing role of information in shaping an agent’s actions, expected returns, and predictions as episodes progress. Additional examples are provided in Appendix B. Unless stated otherwise, all examples are undiscounted and explain optimal policies computed via value iteration (Sutton and Barto, 2018). Steady-state distributions are either computed exactly or accurately estimated from trajectories; expected returns for the outcome characteristic are computed using value iteration. [^1] Code is available at https://github.com/djeb20/sverl.
@@ 7.1 Mastermind
Figure 3: Shapley value attributions for one complete episode of Mastermind, consisting of three states (top to bottom). First column: episode progression, with the optimal action taken in each state shown in green ink—these are not part of the state. Second column: feature contributions to the probability of selecting the optimal action taken (behaviour). Third column: feature contributions to the expected return (outcome). Fourth column: feature contributions to the estimated expected return (prediction). The right-most column on the board displays the position clue; the left-most column displays the misplaced clue. Shapley values are scaled to lie between 1-1- 1 and 111.
In the classic code-breaking game Mastermind, a secret code is selected, and the player attempts to guess it within a limited number of attempts. In our formulation, each code consists of two letters, where each letter can be either $\mathrm{A}$ or $\mathrm{B}$, resulting in four possible secret codes: $\mathrm{AA}$, $\mathrm{AB}$, $\mathrm{BA}$, and $\mathrm{BB}$. After each guess, the player receives feedback in the form of two clues:
@@
  • Position clue: the number of letters in the player’s guess that are present in the secret code and are in the correct position.
@@
  • Misplaced clue: the number of letters in the player’s guess that are present in the secret code but are not in the correct position.
The position clue is computed first, and matched letters are excluded when calculating the misplaced clue. For example, if the secret code is $\mathrm{AB}$ and the guess is $\mathrm{AA}$, the position clue is 1 and the misplaced clue is 0.
We formalise this game as a Markov decision process. At the start of each episode, the environment randomly selects one of the four possible codes as the hidden target. The agent interacts with the environment over three decision stages, submitting one guess per stage. After each guess, the environment returns the position and misplaced clues as observation feedback. The episode terminates when the agent guesses the correct code or after three unsuccessful attempts. Incorrect guesses receive a reward of 1-1- 1, and correct guesses receive 0. Each state is represented by the visible game board, encoded as 16 features consisting of the agent’s previous guesses and their associated feedback. If fewer than three guesses have been made, the remaining feature values are set to a special empty token to indicate the absence of information.
We use SVERL to analyse an agent in Mastermind to show how explanations of behaviour, outcomes, and predictions can align or diverge to reveal complementary insights. Figure 3 shows the state-trajectory for an optimal deterministic policy (left column) and corresponding explanations for an episode where the hidden code is $\mathrm{AB}$.
At the first decision stage, the episode begins on the empty game board. Based on this state, the agent selects $\mathrm{AA}$—though any initial action is optimal. The explanation of this behaviour assigns large positive attributions to the empty bottom row, corresponding to where the agent will place its first code. The prediction explanation assigns large negative attributions to the same features, while the outcome explanation assigns no responsibility to any feature. These patterns reflect that the initial state, indicated by the empty bottom row, has a low value estimate and determines the agent’s initial action. They also reflect that the return is invariant to the known features because all possible initial actions yield the same expected return.
At the second decision stage, the agent can conclude that one letter in its initial guess is correctly placed and the other should be a $\mathrm{B}$. Consequently, it selects $\mathrm{BA}$. The explanations of this behaviour and its outcome assign large positive attributions to the position clue; this clue narrows the set of possible codes. The prediction explanation, in contrast, assigns a large negative attribution to the position clue; a clue of 111 leaves two possible hidden codes, while a clue of 0 would have ruled out all but one. The agent’s expected return is lower under the observed clue than it would be under a more informative one.
@@
In the final state, the agent selects $\mathrm{AB}$ after receiving clues uniquely identifying the hidden code. The features corresponding to the previous guess and its feedback receive high attribution across all three explanations: they narrow down the optimal action (behaviour), which yields the high return received (outcome), and signal that the agent is in a state with maximal expected return (prediction).
@@ 7.2 Minesweeper
Figure 4: Shapley value attributions for two successive states in Minesweeper (top to bottom). First column: episode progression. Second column: feature contributions to the probability of selecting the optimal action taken (behaviour). Third column: contributions to the agent’s expected return (outcome). Fourth column: contributions to the agent’s estimated expected return (prediction). Mines are hidden from the agent; markers “M1” and “M2” are shown for reference. Shapley values are scaled to lie between 1-1- 1 and 111.
Minesweeper is a puzzle in which the player must open squares on a grid while avoiding hidden mines. Selecting a square reveals either a mine (ending the game) or a number indicating how many adjacent squares contain mines. If a square with zero adjacent mines is revealed, its neighbouring squares are recursively opened until nonzero borders are reached. The game ends when either all safe squares have been uncovered or a mine is selected. Our formulation is played on a $\times 4$ 4 grid, and each episode begins with two mines randomly placed and all squares unopened. We model the game as a Markov decision process with approximately 175,000 unique states: the agent selects an unopened square at each decision stage, receiving a reward of 1-1- 1 if it reveals a mine (terminating the episode), and 0 otherwise. Each state is represented by the visible game board, encoded as 16 features—one for each square—taking values unopened, or 0, 1 or 2 indicating the number of adjacent mines.
We use SVERL to analyse an agent in Minesweeper to show how explanations of behaviour, outcomes, and predictions scale to the feasible limits of exact Shapley value computation. We explain a policy trained using Q-learning (Watkins and Dayan, 1992), with the expected return from outcome characteristic function estimated via Monte Carlo rollouts. Figure 4 shows SVERL’s explanations for states from two sequential decision stages.
At the first decision stage, the agent opens square (4,2)(4,2)( 4 , 2 ); at the second, it opens square (4,4)(4,4)( 4 , 4 ). In both stages, the behaviour and outcome explanations assign the highest attribution to the feature at (4,2)(4,2)( 4 , 2 ). Many features help identify the location of the first mine (M1), but only the clue revealed at square (4,2)(4,2)( 4 , 2 ) uniquely determines the location of the second mine (M2). SVERL captures this by assigning the highest attribution to that feature, reflecting its role in determining the agent’s optimal action and the resulting expected return.
The prediction explanation shows a consistent pattern: positive attributions to opened squares and negative attributions to unopened ones. States with more unopened squares often carry greater uncertainty about mine locations and thus often have lower expected returns. SVERL’s explanation of predictions reflects this relationship, where open squares, associated with greater certainty, lead to higher return estimates.